Question:

Suppose that business travelers and vacationers have the following demand for airline tickets from New York to Boston:

PriceQuantity Demanded (business travelers)Quantity Demanded (vacationers) $150 2,100tickets 1,000tickets 200 2,000 800 250 1,900 600 300 1,800 400 (a.) As the price of tickets rises from $200 to $250, what is the price elasticity of demand for (i) business travelers and (ii) vacationers?


1
Expert's answer
2021-10-20T10:49:21-0400

1) Price elasticity of demand for business travellers.

Formula.

Percent change in quantity demanded/percent change in price.

Percent change in quantity demanded "= \frac{q_2-q_1}{q_1}"

"=\frac{2000-1900}{2000} \\

=\frac{100}{2000} \\

=0.05"

Percent change in price "= \frac{p_2-p_1}{p_1}"

"=\frac{200-250}{200} \\

=\frac{-50}{200} \\

=-0.25"

Therefore.

"\frac{0.05}{-0.25}=-0.2"

Therefore price elasticity of demand for business travellers=-0.2

2) For vacationers.

Percent change in quantity demanded

"= \frac{800-600}{800} \\

=\frac{200}{800} \\

=0.25 \\

\frac{0.25}{-0.25} =-1"

Therefore price elasticity of demand for vacationers=-1